What is the first five terms and the explicit formula to a11= 14 and a39= 98
Answer by richwmiller(17219)

a+38d=98
a+10d=14
28d=84
4d=12
d=3
a+30=14
a=-16
-16+114=98
an = -16 + (n – 1)3
an=3n-19
first 5 terms
-16,-13,-10,-7,-4
-1,2,5,8,11, 14 (a11)